26 Feb 2008: Early JET Results
The decision was made to switch the JET from Yellow and Blue every two running days, excluding APEX and Maintenance Day and days of catastrophic running. This gives enough data for a ~5% relative measurement of the polarization each period.
01 Nov 2007: All new detectors installed. Some vacuum problems have prevented the JET from reaching operating vacuum. Preamps boxes have been removed to improve cooling.
27 Sept 2007: We have replacement silicon detectors in hand. They are currently undergoing testing. Two channels initially tested bad, but the trouble was tracked down to wire bonding problems that have been fixed. All channels appear to operate now. Installation is expected in mid October.
Scheduling Physicist: Mike Sivertz | Last updated: |
Email: sivertz@bnl.gov | Office Phone: 631-344-6102 | Digital Pager: 7528 | Home Phone: 631-878-9020 |